OpenStudy (meilendurcer):

How do you simplify this expression? 9 • (7 –4)2 + 9

OpenStudy (anonymous):

(3)^2.(3)^2+3^2 =3^4+3^2

OpenStudy (anonymous):

3^2(3^2+1)=9(9+1)=90

OpenStudy (anonymous):

9 • (7 –4)2 + 9 9 • (3)2 + 9 9 • 6 + 9 36 + 9 45 .....................wait, is the 2 an exponent. If so then it is: 9 • (7 –4)2 + 9 9 • (3)2 + 9 9 • 9 + 9 81 +9 90